Suppose that the Treasury bill rate is 4% and the expected return on the market stays at 9%. Use the following information.

Stock Beta (β)
Caterpillar 1.68
Dow Chemical 1.63
Ford 1.42
Microsoft 0.96
Apple 0.93
Johnson & Johnson 0.55
Walmart 0.47
Campbell Soup 0.37
Consolidated Edison 0.19
Newmont 0.00

a. Calculate the expected return from Johnson & Johnson. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Expected return             %

b. Find the highest expected return that is offered by one of these stocks. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

Highest expected return             %

c. Find the lowest expected return that is offered by one of these stocks. (Do not round intermediate calculations. Enter your answer as a percent rounded to the nearest whole number.)

Lowest expected return             %

d. Would Ford offer a higher or lower expected return if the interest rate were 2% rather than 4%? Assume that the expected market return stays at 9%.

Higher
Lower

e. Would Walmart offer a higher or lower expected return if the interest rate were 8% rather than 4%? Assume the expected market return stays at 9%.

Higher
Lower
Answer #1

1.

=4%+0.55*(9%-4%)=6.75%

2.

=4%+1.68*(9%-4%)=12.40%

3.

=4%+0.00*(9%-4%)=4%

4.

HIGHER

OLD=4%+1.42*(9%-4%)=11.10%

NEW=2%+1.42*(9%-2%)=11.94%

5.

HIGHER

Old=4%+0.47*(9%-4%)=6.35%

New=8%+0.47*(9%-8%)=8.47%
Answer #2

To calculate the expected return for Johnson & Johnson and find the highest and lowest expected returns among the given stocks, we will use the Capital Asset Pricing Model (CAPM):

Expected Return = Risk-free Rate + Beta * (Expected Market Return - Risk-free Rate)

Given information: Treasury bill rate (Risk-free Rate) = 4% Expected market return = 9%

a. Johnson & Johnson: Beta (β) = 0.55

Expected Return = 4% + 0.55 * (9% - 4%) = 4% + 0.55 * 5% = 4% + 2.75% = 6.75%

The expected return for Johnson & Johnson is 6.75%.

b. Highest Expected Return: Among the given stocks, we need to find the one with the highest expected return. By calculating the expected return using the CAPM equation for each stock and comparing them, we can determine the highest expected return.

The highest expected return is for Caterpillar with a beta of 1.68.

Expected Return = 4% + 1.68 * (9% - 4%) = 4% + 1.68 * 5% = 4% + 8.4% = 12.4%

The highest expected return among the given stocks is 12.4%.

c. Lowest Expected Return: Similar to finding the highest expected return, we can determine the lowest expected return among the given stocks by calculating the expected return using the CAPM equation for each stock and comparing them.

The lowest expected return is for Newmont with a beta of 0.00.

Expected Return = 4% + 0.00 * (9% - 4%) = 4% + 0.00 * 5% = 4% + 0% = 4%

The lowest expected return among the given stocks is 4%.

d. Ford and Interest Rate of 2%: If the interest rate were 2% instead of 4%, assuming the expected market return remains at 9%, we can calculate whether Ford would offer a higher or lower expected return.

Ford Beta (β) = 1.42

Expected Return (Interest Rate = 2%) = 2% + 1.42 * (9% - 2%) = 2% + 1.42 * 7% = 2% + 9.94% = 11.94%

If the interest rate were 2% instead of 4%, Ford would offer a higher expected return of 11.94% compared to the previous expected return of 4% + 1.42 * 5% = 11.1%.

e. Walmart and Interest Rate of 8%: If the interest rate were 8% instead of 4%, assuming the expected market return remains at 9%, we can calculate whether Walmart would offer a higher or lower expected return.

Walmart Beta (β) = 0.47

Expected Return (Interest Rate = 8%) = 8% + 0.47 * (9% - 8%) = 8% + 0.47 * 1% = 8% + 0.47% = 8.47%

If the interest rate were 8% instead of 4%, Walmart would offer a higher expected return of 8.47% compared to the previous expected return of 4% + 0.47 * 5% = 6.35%.


answered by: mervetokaz